Find the mass of ammonia in the 100 cubic feet tank having a pressure of 70 psi at 120 °F.


1
Expert's answer
2021-10-10T15:59:15-0400

Ideal gas law

"PV=nRT \\

V=100 \;ft^3 = 2831.68 \;L\\

P=70 \;psi = 4.7632 \;atm\\

T= 120 \;°F = 322.039"

R=0.082058 L×atm/(K×mol)

"n = \frac{PV}{RT} \\

= \frac{4.7632 \times 2831.68}{0.082058 \times 322.039} \\

= 510.40 \;mol"

M(ammonia) = 17.031 g/mol

m(ammonia) "= 510.40 \times 17.031 = 8692.62 \;g = 8.692 \;kg"

Answer: 8.7 kg